DISCOVERY OF MASSIVE YOUNG STELLAR OBJECTS IN THE GALACTIC CENTER WITH WARM CO_{2} GAS ABSORPTION

DEOKKEUN AN, SOLANGE RAMIREZ, IPAC/CALTECH; \underline{KRIS SELLGREN}, OHIO STATE U.; ADWIN BOOGERT, CALTECH; RICHARD ARENDT, NASA/GODDARD SPACE FLIGHT CENTER; ANGELA COTERA, SETI INSTITUTE; THOMAS ROBITAILLE, ST. ANDREWS U., UK; MATHIAS SCHULTHEIS, OBS. BESANCON, FRANCE; HOWARD A. SMITH, HARVARD-SMITHSONIAN CENTER FOR ASTROPHYSICS; SUSAN STOLOVY, SPITZER SCIENCE CEN-TER, CALTECH.

We report the detection of several molecular gas-phase and ice absorption features in three photometrically-selected young stellar object (YSOs) candidates in the central 280 pc of the Milky Way. Our spectra, obtained with the Infrared Spectrograph (IRS) onboard the Spitzer Space Telescope, reveal gas-phase absorption from CO_{2} (15.0 μm), C_{2}H_{2} (13.7 μm) and HCN (14.0 μm). We attribute this absorption to warm, dense gas in the massive YSOs. We also detect strong and broad 15.2 μm CO_{2} ice absorption features, with an absorption profile indicating the presence of thermally processed CO_{2} ice. Our IRS observations demonstrate the youth of these objects, and provide the first spectroscopic identification of massive YSOs in the Galactic Center.